John can type 2/3 manuscript in 12 hours. With Laura's help they can complete the typing in 2 hours. How many hours would Laura take to type the manuscript working alone?
Answer by nerdybill(7384)

John can type 2/3 manuscript in 12 hours. With Laura's help they can complete the typing in 2 hours. How many hours would Laura take to type the manuscript working alone?
.
Let x = rate of Laura's (manuscript per hour)
.
time * rate = amount of manuscript completed
therefore, "manuscript completed by John in 2 hrs" + "manuscripts completed by Laura in 2 hours" = one manuscript
.
(2)[(2/3)/12] + 2x = 1
(2/3)/6 + 2x = 1
2/18 + 2x = 1
2 + 36x = 18
36x = 16
x = 16/36
x = 4/9 (four manuscript per 9 hours)
.
For one manuscript:
Let t = time
4t/9 = 1
4t = 9
t = 9/4 of an hour
t = 2 1/4 hours
or
t = 2 hours 15 minutes